OpenStudy (thatonegirl_):

more log questions.. 4e^(-2x)=17 solve for x?

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

Use slader btw, if its book work. Also, wolframalpha if yuo don't care for work now 4e^(-2x)=17 e^(-2x)=17/4 ln(e^(-2x))=ln(17/4) (-2x)ln(e)=ln(17/4) -2x=ln(17/4) x=(ln(17/4))/(-2)

OpenStudy (thatonegirl_):

thanks!